201111 R1ProgressRep

From New IAC Wiki
Jump to navigation Jump to search


70 wires were strung today. Our second stringer should be starting tomorrow.


91 wires were strung today.

20111102 Temp1.png


The correct HVAC filters came in today. A clean filter was installed today at ~10 AM. A small rise in temperature can be seen in the temperature graph when this change was made.

77 wires were strung today.

20111103 Temp1.png


91 wires were strung today.

All field wires are strung on SL2.

20111104 Wires.png

20111104 Temp2.png


Spent the day visually inspecting for crossed wires. We should have the chamber rotated tomorrow and hopefully back to stringing later in the afternoon.

20111107 Temp1.png


Chamber was rotated this morning.

We started stringing SL1 mid-afternoon. We only managed 50 wires today.

20111108 Temp1.png

Temperature drop was caused by our power supply overheating. For future reference, don't exceed the rated 60 Amps of current...


I removed one of the covers for our large ducting in an attempt to get some air flow through the room (not the clean room). This appears to have changed the cyclic rate of our HVAC. Temperature min/max appears to have been unaffected.

20111109 Temp1.png

The ducting cover will be put back in place tomorrow.

63 wires were strung today.

20111109 Temp2.png


63 wires were strung today.

Keeping the duct cover off until tomorrow.


The duct cover was put back on about 1 PM.

20111111 Temp1.png

105 wires were strung today.

20111111 Temp2.png


126 wires were strung today.

20111114 Temp1.png


119 wires were strung today. We also changed our spool of field wire.

20111115 Temp1.png


126 wires were strung today.

20111116 Temp1.png


126 wires were strung today. We are at about the halfway point for SL1 (row 112).

20111117 Temp1.png


Took the girls shoe shopping this afternoon. Cleanroom shoes are on order now.

We still managed to do 126 wires today.


I came in to do our weekly trend measurements. Power supply was not warm to the touch, so not sure what was going on with temperatures.

20111119 Temp1.png


112 wires were strung today.

20111121 Temp1.png


126 wires were strung today.

20111122 Temp1.png


I spent some time this afternoon testing and adjusting our HVAC temperature alarm. The alarm now goes off if temperatures exceed 72° F and if it falls below 68° F inside the cleanroom.


56 wires were strung today.

20111128 Temp1.png


105 wires were strung today.


84 wires were strung today.

400 px